A company’s second year operations can be summarized as follows:
Revenues: $110 000
Expenses (except CCA): $65 000
Its capital asset purchases in the first year totalled $100 000, with none in the second year. With a CCA rate of 20 percent and a tax rate of 55 percent, how much income tax did it pay?
1
Expert's answer
2020-05-25T09:19:15-0400

"Net income = 110000-65000-50000(0.2)="

"45000-10000=35000"

"Therefore taxes paid = 35000(0.55)=\$19250"